# Alchemical harmonic approximation based potential for all iso-electronic diatomics: Foundational baseline for $\Delta$-machine learning

### v1

@@ Abstract
We introduce the alchemical harmonic approximation (AHA) model of the absolute electronic energy for all stoichiometries corresponding to charge-neutral iso-electronic diatomics at some fixed interatomic distance $d_{0}$. To account for variations in interatomic distance, we combine AHA with the following Ansatz for the electronic contribution to the binding potential, $E(d)=(E_{u}-E_{s})\left(\frac{E_{c}-E_{s}}{E_{u}-E_{s}}\right)^{\sqrt{d/d_{0}}}+E_{s}$, where $E_{u}$, $E_{c}$, and $E_{s}$ correspond to the energies of the united atom, calibration at $d_{0}$, and sum of infinitely separated atoms, respectively. For any number of electrons, our model covers the entire two-dimensional electronic potential energy surface spanned by interatomic distance and difference in nuclear charge from which only one single point (fixed elements of nuclear charge $Z_{1},Z_{2}$ and distance $d_{0}$) is drawn to calibrate ($E_{c}$). Using data from pbe0/ccpvdz as a reference, we present numerical evidence for the electronic ground-state potentials of all neutral diatomics with 8, 10, 12, 14 electrons. We assess the validity of our model by comparison to legacy interatomic potentials (Harmonic oscillator, Lennard-Jones, and Morse) within the most relevant range of binding (0.7 - 2.5 Å), and find comparable accuracy if restricted to single diatomics, and significantly better (and more systematic) predictive power when extrapolating to the entire iso-electronic series. We have also investigated $\Delta$-learning of the electronic absolute energy for any iso-electronic diatomic, using our model as a baseline. For any given iso-electronic charge neutral diatomic surface, our baseline model results in a systematic improvement, effectively reducing training data needs for reaching chemical accuracy ($\sim$ kcal/mol) by up to an order of magnitude from $\sim 1000$ to $\sim 100$ training instances. By contrast and with respect to direct learning, using AHA+Morse as a baseline hardly leads to any improvement, and sometimes even deteriorates the predictive power. Direct KRR-based extrapolation throughout chemical space converges to a prediction error of $\sim$0.1 Ha when inferring the energy of unseen CO after training on all other iso-electronic diatomics. Inclusion of our model as a baseline (calibrated to the energy of BF at $d_{0}=1.2$) lowers the error to $\sim$0.04 Ha.
@@ I Introduction
Quantum mechanics underpins our ability to predict electronic, optical, and thermal properties with high fidelity, essential for understanding chemical space or designing materials with specific functionalities ceder1998predicting . Unfortunately, numerically solving meaningful approximations to the electronic Schrödinger equation on-the-fly for each and every material remains a computationally prohibitive challenge. Significant acceleration can be achieved via machine learning (ML) based inference which can remove the need for its solution rupp_muller_lilienfeld_2012 ; Elpasolite_2016 ; QMLessayAnatole , or enhance the numerical methods apbe0 commonly employed. Although practical and more universally applicable by now, e.g. in the form of fragment based building blocks for quantum machine learning Amons ; huang2021abinitio , or "general-purpose" models of force field potentials trained on diverse data batatia2024foundationmodelatomisticmaterials , freely sampling chemical space is still hampered due to ML models’ inherent interpolative nature which lacks the universality of the Schrödinger equation. Quantum alchemy based techniques by contrast provide an interesting alternative, physically motivated scheme that exploits perturbation theory to bypass explicit solutions across different composition, i.e. for new systems rudorff2020_alchem_chirality ; krug_generalAIT . In contrast to machine learning, typically only one or very few solutions are required for calibration. While less commonly deployed than their ML counterpart, quantum alchemical perturbation von_Rudorff_2020 based methods have become easily applicable to general systems, e.g. Refs. rudorff2021arbitrarily, ; lilienfeld_tuckerman, ; weigend_firstalchemy, ; michael_hammett_JACS, ; balawender_2019, ; Alchemy_bindingenergies, ; williams-noonan_2018, ; Samuel2018bandgaps, . Less attention has been paid to combining alchemical and geometrical changes, first introduced recently in Ref. domenichini_2022_al_geom_relax, , which is what this work focuses on: an alchemical approximation to nuclear transmutations of diatomics, together with an interatomic potential for different distances $d$, and both as a baseline for a subsequent ML model. Finding such models, i.e. quantitative descriptions of the interatomic potential for different nuclear charges, is an ongoing quest in different energies regimes, e.g. for van der Waals bondsmatej_2024_QDO . As many-body dispersion interactions are often harder to capture than covalent bonds, we will focus on the latter in this work.
@@ II.1 Alchemy
Previous work has shown the convergence behavior of this seriesrudorff2021arbitrarily , as well as its satisfactory accuracy, not just for symmetric systems like diatomics where every odd-numbered term of Eq. 1 vanishesvonlilienfeld2023_1st_order ; rudorff2020_alchem_chirality , but for truncation after its second orderfias_2019 . Then, given such a parabola in $\lambda$, we seek suitable conditions to determine its three degrees of freedom. We are not bound to truncate Eq. 1 after the second order, but testing revealed numerical problems when considering higher orders.
@@ II.2 The alchemical harmonic approximation (AHA)
To extend the energy prediction of the AHA from mere iso-electronic changes in nuclear composition to changes in distance $d$, we seek to construct the parabola of the AHA such that only one of the three necessary constraints exhibits $d$-dependency. For this, pick the endpoints of the energy parabola (where $\lambda=\pm(Z_{1}+Z_{2})/2=:\pm\lambda_{u}$), i.e. the two united atoms ($Z_{u}:=Z_{1}+Z_{2}$) with energy $E_{u}$. The united atoms are obviously independent of any interatomic distance. The third, remaining constraint is given by a point in between, $E(\lambda_{0},d)$, which will change with varying distance. Our calibration point can be found at $d_{0}$, i.e. $E_{0}(d_{0})=E_{c}$.
@@ II.3 The Alchemical Quartic Approximation (AQuA)
Another example is the inclusion of the united atom at only one side of the parbaola:
@@ II.4 Distance-dependence
There are two reasons to consider only the electronic distance-dependence: firstly, alchemy only makes relative statements about the electronic part of Schrödinger equations and secondly, as the Coulomb repulsion is known analytically, it appears only natural to model electronic degrees of freedom separately. However, while Eq. 18 exhibits the correct behavior within the equilibrium range of binding (0.7 - 2.5 Å), and at infinite distance, the total energy upon addition of Coulomb repulsion does not show yet the desired physical dissociation behavior in the long range. In order to also treat dissociation, our potential would still need to be morphed into the correct attractive Coulombic and van der Waals dispersion terms which are well known HeitlerLondon ; EisenschitzLondon , and continue to be further improved and developed matej_2024_QDO . Furthermore, difficulties are to be expected as the regime of larger interatomic distances is affected by electronic multi-reference effects. A viualization of different (electronic, nuclear and total) energies of different diatomics and different distances $d$ can be found in Fig. 2. It serves as complement to Fig. 1 such that slices of one graph constitute the functions of the other.
@@ III.2 Machine Learning with AHA as baseline
Figure 5: Learning curves of Kernel Ridge Regression (KRR) with calibration system BF. ML models correspond to no baseline model (direct learning, dashed line), with our potential and AHA ($\Delta$-learning, solid line) and the Morse potential and AHA ($\Delta$-learning, dotted line). Exponents $n$ of CM(n) are annotated. Test and training set drawn at random. Exponents $n$ of CM(n) are annotated. Inset: Test set exclusively contains CO-samples while training samples are drawn at random from all systems but CO.
[…]
In addition, this procedure can be repeated for different calibration systems (N2, CO, BeNe, LiNa, HeMg, HAl) and the iso-electronic series’ of 8 (Be2, LiB, HeC, HN), 10 (B2, BeC, LiN, HeO, HF) and 12 (C2, BN, BeO, LiF, HeNe, HNa) electrons, with similar results (Fig. 6). Odd numbers of electrons are possible as well, but lead to half-integer $\lambda_{m},\lambda_{0},\lambda_{u}$ when considering physical diatomics.
Clearly, the small, global representation CM(n) works best albeit only marginally. This small difference between global and local kernels diminishes the smaller the parabola of the AHA becomes (N2 to C2 to B2 to Be2). This is to be expected since diatomics do not possess three- or higher-order many-body terms and can be adequately described using the interatomic distance information.
@@ IV Conclusion
Future extensions of this research might deal with (i) inclusion of the interatomic distance dependent alchemical force $F_{0}$ into AHA based models, (ii) extensions to molecules via fragment (amon) based $\Delta$-machine learning Amons ; delta2015 , and (iii) the extension of AHA-baseline $\Delta$-learning to multi-level learning zaspel2018boosting ; heinen2024reducing .
@@ Acknowledgements
We acknowledge discussions with Florian Bley and Oliver Eberle. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2023-04853]. Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2023-04853]. This research was undertaken thanks in part to funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant number: CFREF-2022-00042. O.A.v.L. has received support as the Ed Clark Chair of Advanced Materials and as a Canada CIFAR AI Chair. O.A.v.L. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772834).

### v2

@@ Abstract
We introduce the alchemical harmonic approximation (AHA) of the absolute electronic energy for charge-neutral iso-electronic diatomics at fixed interatomic distance $d_{0}$. To account for variations in distance, we combine AHA with this Ansatz for the electronic binding potential, $E(d)=(E_{u}-E_{s})\left(\frac{E_{c}-E_{s}}{E_{u}-E_{s}}\right)^{\sqrt{d/d_{0}}}+E_{s}$, where $E_{u},E_{c},E_{s}$ correspond to the energies of united atom, calibration at $d_{0}$, and sum of infinitely separated atoms, respectively. Our model covers the entire two-dimensional electronic potential energy surface spanned by distance and difference in nuclear charge from which only one single point (with elements of nuclear charge $Z_{1},Z_{2}$ and distance $d_{0}$) is drawn to calibrate $E_{c}$. Using reference data from pbe0/cc-pVDZ, we present numerical evidence for the electronic ground-state of all neutral diatomics with 8, 10, 12, 14 electrons. We assess the validity of our model by comparison to legacy interatomic potentials (Harmonic oscillator, Lennard-Jones, and Morse) within the most relevant range of binding (0.7-2.5 Å), and find comparable accuracy if restricted to single diatomics, and significantly better predictive power when extrapolating to the entire iso-electronic series. We also investigated $\Delta$-learning of the electronic absolute energy using our model as baseline. This baseline model results in a systematic improvement, effectively reducing training data needs for reaching chemical accuracy by up to an order of magnitude from $\sim$1000 to $\sim$100. By contrast, using AHA+Morse as a baseline hardly leads to any improvement, and sometimes even deteriorates the predictive power. Inferring the energy of unseen CO converges to a prediction error of $\sim$0.1 Ha in direct learning, and $\sim$0.04 Ha with our baseline.
@@ I Introduction
Quantum mechanics underpins our ability to predict electronic, optical, and thermal properties with high fidelity, essential for understanding chemical space or designing materials with specific functionalities ceder1998predicting . Unfortunately, numerically solving meaningful approximations to the electronic Schrödinger equation on-the-fly for each and every material remains a computationally prohibitive challenge. Significant acceleration can be achieved via machine learning (ML) based inference which can remove the need for its solution rupp_muller_lilienfeld_2012 ; Elpasolite_2016 ; QMLessayAnatole , or enhance the numerical methods apbe0 commonly employed. Although practical and more universally applicable by now, e.g. in the form of fragment based building blocks for quantum machine learning Amons ; huang2021abinitio , or "general-purpose" models of force field potentials trained on diverse data batatia2024foundationmodelatomisticmaterials , freely sampling chemical space is still hampered due to ML models’ inherent interpolative nature which lacks the universality of the Schrödinger equation. Quantum alchemy based techniques by contrast provide an interesting alternative, physically motivated scheme that exploits perturbation theory to bypass explicit solutions across different composition, i.e. for new systems AlchemyAlisa_2016 ; Samuel-JCP2016 ; Yasmine-JCP2017 ; Samuel2018bandgaps ; rudorff2020_alchem_chirality ; shiraogawa2023optimization ; krug_generalAIT . In contrast to machine learning, typically only one or very few solutions are required for calibration. While less commonly deployed than their ML counterpart, quantum alchemical perturbation von_Rudorff_2020 based methods have become universally applicable across chemical space, e.g. Refs. lilienfeld_tuckerman, ; weigend_firstalchemy, ; michael_hammett_JACS, ; balawender_2019, ; Alchemy_bindingenergies, ; williams-noonan_2018, ; Samuel2018bandgaps, ; rudorff2021arbitrarily, , and have even already been used to calculate meaningful baselines for $\Delta$-machine learning models of catalytic activity griego2020machine .
In this work we focus on alchemically motivated approximate interpolations of the entire series of nuclear transmutations in diatomics possible for a given number of electrons. Motivated by the concavity of the electronic potential in any alchemical changes anatole-ijqc2013 , we have investigated low order even polynomials, i.e. the alchemical harmonic approximation (AHA) and the alchemical quartic approximation (AQuA). In order to also account for the impact of interatomic distances, a novel potential has been developed. Due to its interpolative nature, our joint universal potential model of alchemical and geometrical changes goes beyond our recent preceding efforts to estimate geometry changes due to compositional changes domenichini_2022_al_geom_relax via perturbation theory and using the mixed Hessian involving nuclear charges and atomic positions (‘alchemical forces’) fias_2019 . Finding such models, i.e. quantitative descriptions of the interatomic potential for different nuclear charges, is an ongoing quest in different energies regimes, e.g. for van der Waals bondskhabibrakhmanov2023universal ; matej_2024_QDO . As many-body dispersion interactions are typically dominating the interatomic long-range regime, this work focuses on the short-range distances governed by covalent binding.
@@ II.1 Alchemy
Previous work has shown the convergence behavior of this seriesrudorff2021arbitrarily , as well as its satisfactory accuracy, not just for symmetric systems like diatomics where every odd-numbered term of Eq. 1 vanishesvonlilienfeld2023_1st_order ; rudorff2020_alchem_chirality , but for truncation after its second orderfias_2019 .
Given the strict condition that the $E$ is concave in any $\lambda$ anatole-ijqc2013 , we begin by assuming the shape of a parabola, and we seek suitable conditions to determine its three degrees of freedom. We are not bound to truncate Eq. 1 after the second order, but testing revealed numerical instabilities when considering higher orders. Also note that this quadratic Ansatz attempts to directly model the expectation value of the electronic Schrödinger equation, in line with our recently published model of the energy of the free atom krug2024energies ; and is to be contrasted with previous attempts to replace alchemical linear interpolations of the Hamiltonian by quadratic interpolations accurate_abinitio .
@@ II.2 The alchemical harmonic approximation (AHA)
To extend the energy prediction of the AHA from mere iso-electronic changes in nuclear composition to changes in distance $d$, we seek to construct the parabola of the AHA such that only one of the three necessary constraints exhibits $d$-dependency. For this, pick the endpoints of the energy parabola (where $\lambda=\pm(Z_{1}+Z_{2})/2=:\pm\lambda_{u}$), i.e. the two united atoms ($Z_{u}:=Z_{1}+Z_{2}$) with energy $E_{u}$ which is independent of interatomic distance. The third, remaining constraint is given by a point in between, $E(\lambda_{0},d)$, which will change with varying distance. Our calibration point can be found at $d_{0}$, i.e. $E_{0}(d_{0})=E_{c}$.
@@
Note, that both $E_{u}$ and $E_{c}$ must be in the same electronic state, e.g. computing $E_{c}$ at a small distance via quantum chemistry software will prove difficult as most methods implicitly assume the Born-Oppenheimer approximation, neglecting nuclear dynamics. Consequently, care must be taken, both with the calibration distance $d_{0}$ and the corresponding electronic state.
@@ II.3 The Alchemical Quartic Approximation (AQuA)
Another example is the inclusion of the united atom at only one side of the parabola:
@@ II.4 Distance-dependence
There are two reasons to consider only the electronic distance-dependence: firstly, alchemy only makes relative statements about the electronic part of Schrödinger equations and secondly, as the Coulomb repulsion is known analytically, it appears only natural to model electronic degrees of freedom separately. However, while Eq. 18 exhibits the correct behavior within the equilibrium range of binding (0.7 - 2.5 Å), and at infinite distance, the total energy upon addition of Coulomb repulsion does not show yet the desired physical dissociation behavior in the long range. In order to also treat dissociation, our potential would still need to be morphed into the correct attractive Coulombic and van der Waals dispersion terms which are well known HeitlerLondon ; EisenschitzLondon , and continue to be further improved and developed khabibrakhmanov2023universal ; matej_2024_QDO . Furthermore, difficulties are to be expected as the regime of larger interatomic distances is affected by electronic multi-reference effects. A viualization of different (electronic, nuclear and total) energies of different diatomics and different distances $d$ can be found in Fig. 2. It serves as complement to Fig. 1 such that slices of one graph constitute the functions of the other.
@@ III.2 Machine Learning with AHA as baseline
Figure 5: Learning curves of Kernel Ridge Regression (KRR) with calibration system BF. ML models correspond to no baseline model (direct learning, dashed line), with our potential and AHA ($\Delta$-learning, solid line) and the Morse potential and AHA ($\Delta$-learning, dotted line). Exponents $n$ of CM(n) are annotated. Test and training set drawn at random. Inset: Test set exclusively contains CO-samples while training samples are drawn at random from all systems but CO.
As a calibration calculation, we again pick BF. Then, the difference between ours and the Morse potential as a baseline model for the prediction of diatomic energies and the reference data at DFT-level of theory (pbe0/cc-pVDZ) can be learned with KRR. For this, we generate 7 times 1024 points (for the 7 different $\lambda$ and 1024 points for $d\in[0.7,2.5]$   Å), then choose different global and local representations for the diatomics (see Sec. Computational Details). The kernel is Laplacian with the Manhattan norm. To determine hyperparameters, we employ 8-fold cross-validation for different training set sizes $N_{\text{train}}=8,16,32,64\dots,4096$. The size of the test set is $N_{\text{test}}=3072$. We compare these two $\Delta$-learning approaches with direct learning of the CM(n) representation (Fig. 5). Note the different sizes of representations: while CM(n) contains only four numbers per diatomic, cMBDF consists of 40 (times 2 atoms), FCHL19 of 3952 (times 2 atoms) and finally SOAP with 44296 (times 2 atoms)! Especially SOAP scales costly as every element species of the 14-electron diatomics is included, i.e. elements H to Si.
In addition, this procedure can be repeated for different calibration systems (N2, CO, BeNe, LiNa, HeMg, HAl) and the iso-electronic series’ of 8 (Be2, LiB, HeC, HN), 10 (B2, BeC, LiN, HeO, HF) and 12 (C2, BN, BeO, LiF, HeNe, HNa) electrons, with similar results (Fig. 6). Odd numbers of electrons are possible as well, but lead to half-integer $\lambda_{m},\lambda_{0},\lambda_{u}$ when considering physical diatomics. However, we did not study these diatomics yet as it would require more sophisticated treatments of the open-shell electron system.
In Fig. 5, the small, global representation CM(n) clearly works best albeit only marginally. This small difference between global and local kernels diminishes the smaller the parabola of the AHA becomes (N2 to C2 to B2 to Be2, Fig. 6). This is to be expected since diatomics do not possess three- or higher-order many-body terms and can be adequately described using the interatomic distance information.
@@ IV Conclusion
Future extensions of this research might deal with (i) inclusion of the interatomic distance dependent alchemical force $F_{0}$ into AHA based models, (ii) the extension of chemical space by considering chemical species which are not iso-electronic in total electron count but rather in number of valence electrons, as already examplified for ionic crystals AlchemyAlisa_2016 , covalent bonding Samuel-JCP2016 , and band-gaps of (III)-(V) semi-conductors Samuel2018bandgaps , (iii) extensions to molecules via fragment (amon) based $\Delta$-machine learning Amons ; delta2015 , and (iv) the extension of AHA-baseline $\Delta$-learning to multi-level learning zaspel2018boosting ; heinen2024reducing .
@@ Acknowledgements
We acknowledge discussions with Florian Bley, Oliver Eberle and Katharine Hunt. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2023-04853]. Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2023-04853]. This research was undertaken thanks in part to funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant number: CFREF-2022-00042. O.A.v.L. has received support as the Ed Clark Chair of Advanced Materials and as a Canada CIFAR AI Chair. O.A.v.L. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 772834).